# Spread Your Wings: A Radial Strip Transformer for Image Deblurring

### v1

@@ Abstract
Exploring motion information is important for the motion deblurring task. Recent the window-based transformer approaches have achieved decent performance in image deblurring. Note that the motion causing blurry results is usually composed of translation and rotation movements and the window-shift operation in the Cartesian coordinate system by the window-based transformer approaches only directly explores translation motion in orthogonal directions. Thus, these methods have the limitation of modeling the rotation part. To alleviate this problem, we introduce the polar coordinate-based transformer, which has the angles and distance to explore rotation motion and translation information together. In this paper, we propose a Radial Strip Transformer (RST), which is a transformer-based architecture that restores the blur images in a polar coordinate system instead of a Cartesian one. RST contains a dynamic radial embedding module (DRE) to extract the shallow feature by a radial deformable convolution. We design a polar mask layer to generate the offsets for the deformable convolution, which can reshape the convolution kernel along the radius to better capture the rotation motion information. Furthermore, we proposed a radial strip attention solver (RSAS) as deep feature extraction, where the relationship of windows is organized by azimuth and radius. This attention module contains radial strip windows to reweight image features in the polar coordinate, which preserves more useful information in rotation and translation motion together for better recovering the sharp images. Experimental results on six synthesis and real-world datasets prove that our method performs favorably against other SOTA methods for the image deblurring task. The code are available in the https://github.com/Calvin11311/RST.
@@ 1 Introduction
Image deblurring aims to recover a sharp image from the blurred one. This is a challenging problem as both blur and latent clear images are unknown. This problem has been attracted a lot by researchers from both academia and industry due to the rapid development of kinds of smartphones. Significant progress has been made due to kinds of deep convolution neural networks (CNNs) [42, 26, 12]. The key success of these methods is because of the capability to model the local blur pattern by a convolution kernel and the diversity design of network architecture (e.g., the multi-scale [42, 44], multi-stage [26, 36] generative adversarial learning [17, 18], and latent space kernel estimation [12]). Since the limited receptive field and the spatially invariant of the convolution kernels [41] can not fully explore non-local contextual information that is vital for image deblurring, a deeper and larger network may relieve this problem. However, simply increasing the depth or capacity of deep networks does not always lead to better performance [45, 29].
@@
Figure 1: (a) Illustration of the motion field from the kernel estimation method [2], the red curves describe the motion state which causes the blurry images. (b) The demonstration of the motion trajectory causing blurry results can be composed of translation and rotation parts, as the blue curve consists of the red rotation motion and the yellow translation motion. (c) Illustrate the token relationship modeled under the polar coordinate system, where the $\theta_{ab}$ can represent the angular relative position between tokens A and B.
@@
To overcome this problem, several approaches [41, 37, 16] explore the vision Transformer [10] to solve image deblurring and achieve decent performance on kinds of benchmark datasets. However, the scaled dot-product attention in the Transformer is time-consuming. Some methods adopt the rectangle window [6] or the strip windows [37] to compute self-attention on lower computation complexity. Since the restricted spatial extent of self-attention results in limitations for capturing long-range pixel relationships, the window shift operation is necessary to expand the receptive field. The shift operation under the Cartesian coordinate system captures the relationship between windows in orthogonal directions. However, in Figure 1 (a), we can clearly find that the blur kernel in real-world blurry images is an irregular curve indicating that the motion causing blurry images is complex. Furthermore, the motion usually can be decomposed into rotation and translation motion. Deng et al. [31] formulate the decomposition as $RT(o,\theta,t,x)=R(\theta)(x-o)+o+t$, where t and $\theta$ represent the translation vector and the rotation angle, respectively, o and x are the coordinates in the blurry image and the $R(\theta)$ is a geometric rotation matrix. Based on the above theory, as illustrated in Figure 1 (b), the trajectory of motion can consist of the rotation (the red curve) and the translation (the yellow line). We note that attention operations capturing both translation and rotation motion information together can enhance the ability to restore clearer images.
@@
However, shift operation under the Cartesian coordinate system (e.g., Stripformer [37]) only considers the cross-window connection in orthogonal directions, which explicitly models translation motion, while implicitly modeling rotation motion. Resulting in the limited capability to model the rotation information for restoring sharp images. In the polar coordinate system, we can use angles and distance to demonstrate the position relative, where the translation and the rotation can be considered together. So it is more suitable to develop an effective and efficient polar system-based transformer architecture to better explore the motion information of blurry images. In this paper, we propose a Radial-Strip Transformer (RST), exploiting two polar coordinate system-based modules: the dynamic radial embedding (DRE) and radial strip attention solver (RSAS), to capture the rotation and the translation features of motion efficiently. The DRE utilizes a polar mask to construct the shallow feature extractor, which is combined with the deformable convolution to capture the motion information of blur on the polar way. In detail, the rotation information can be captured by angles and the translation part is modeled by distance.
@@
Since features extracted from DRE are not fully utilized by attention operations with the Cartesian coordinate system, directly using window-based transformer architecture (e.g., Swin Transformer [24]) will not result in a good performance of image deblurring. Inspired by the DarSwin [1], which proposes a polar transformer under a wide-angle lens, we design an efficient radial strip attention solver (RSAS), which contains strip windows along the radius. Sampling operations are needed in sector window partitioning but can result in the loss of details, which is important for image deblurring. So we cancel the sampling operation and set the strip window, which is beneficial for preserving the useful detailed information for clear image restoration. Combined with the two designs above, we proposed our RST with an asymmetric encoder-decoder architecture, which only uses the RSAS in the decoder part. It is more efficient to solve the deblurring task because the deep layer is clearer than the shallow ones [16]. The experimental results support that the method performs a favorable deblurring result against SOTA methods.
@@
    We propose an efficient polar coordinate system-based embedding module DRE for extracting rotation and translation motion information of shallow features. It transforms the deformable convolution layer in each radial direction and preserves the information in rotation and translantion.
@@ 2.1 Deep CNN-based approaches
Recently, the development of deep CNN methods [4, 7, 13] has achieved great performance in the image deblurring task. Nah et al. [26] design a muti-scale network to estimate the blur pattern from a blurry image directly. To make the multi-scale architecture extract the blur pattern more efficiently, Gao et al. [13] propose a parameter selective sharing method. [35] utilized the CNNs to predict the probabilistic distribution of the blurring kernel at the patch level. Moreover, the multi-patch architecture is another type to improve the performance of deblurring. Zhang et al. [44] performs a hierarchical deblurring model based on the multi-patch strategy, which estimates the feature of the blur pattern stage by stage. The cross-stage strategy is utilized by Zamir et al. [42], which fuses the different stage features to better restore the sharp image.
Since the above methods only directly use the fixed convolution kernel to estimate the blur pattern, which ignores the important movement prior of the blurry images. Some methods concern the direction of degradation, and WDNet [23] designs a dual-branch network for demoiréing, which contains a DPM block to capture the texture from eight directions. In DSC [14] utilizes the attention mechanism in a spatial recurrent neural network for locating shadows from each direction. Fang et al. [11] tries to estimate the non-uniform blur kernel in latent space and use it to reconstruct the sharp image. Due to the convolution operation being a spatial invariant and limited receptive field [41], CNN methods do not efficiently capture the global information for restoring blurry images.
@@ 2.2 Vision Transformer-based approaches
Transformer architecture achieves a significant performance of modeling global contexts with self-attention in high-level tasks (e.g., moving object segmentation [30], object detection [17], and text recognition [19]), and extend to low-level tasks(image deblurring [16, 41], super-resolution [21], and image denoise [3, 39]). But, self-attention brings a problem of huge computational cost. To solve it, SwinIR [21] divides the image by 8$\times$8 windows and executes an attention operation in each window. Restormer [41] designs across channel dimensions model to reduce the cost. The spatial information is important for the image restoration task, which is not fully developed. CAT [6] aggregates the features across rectangle windows with axial-shift operation, which can achieve a large receptive field and boost the interactions between windows with linear complexity. Stripformer [37] proposed an intra-strip and inter-strip attention mechanism, which decomposes motion information into horizontal and vertical directions for estimating the blur pattern. However, the above methods designed based on the Cartesian coordinate system can only explore the motion information in orthogonal directions. Since translation and rotation are both important parts of object movement, the Cartesian coordinate has a limited ability to model the rotation. Different from vanilla transformers calculate the relationship of pixels on the Cartesian system, we propose an efficient polar coordinate transformer model to explore the spacial information of movement with translation and rotation.
@@ 3.1 Overall Pipeline
Given a blurry image $\bm{I}$ with a resolution of $H\times W$. RST first adopts a deformable convolution layer with a radial mask to extract the shallow feature $\bm{F}_{0}\in\mathbb{R}^{H\times W\times C}$, where $H,~{}W,~{}C$ denote the height, width, and number of channels, respectively. Then the shallow feature $\bm{F}_{0}$ is embedded in the L-level encoder-decoder networks where each level has a skip connection between the encoder and decoder. Following the pioneered work [16], we directly use the frequency domain-based feed-forward network to refine the features, which has the benefit of capturing useful context information for deblurring. The process can be formulated as:
@@
$\mathcal{F}(\cdot)$ denotes the fast Fourier transform and $\mathcal{F}^{-1}(\cdot)$ denotes the inverse one. $\bm{W}$ represents a learnable quantization matrix. $\bm{F}_{e}\in\mathbb{R}^{H\times W\times C}$ represents the output feature from FFN of encoder parts. We use the transpose convolution for downsampling and upsampling and add the skip connection between the encoder and decoder on each level. In the decoder, we design radial strip attention (as shown in Figure 2) to extract the blur pattern from the rotation and translation movement. To be specific, the feature $\bm{F}_{a}\in\mathbb{R}^{\frac{H}{M}\times\frac{W}{M}\times C}$ ($M$ represents the patch size) product by window partition layer. After that, the windows are assembled by an azimuth patch merging module with the angular relative positional encoding. The transformer block outputs the final restored feature $\bm{F}_{r}\in\mathbb{R}^{H\times W\times C}$, and using a $3\times 3$ convolution layer to generate the residual part $\bm{R}\in\mathbb{R}^{3\times H\times W}$. Finally, the restored image is generated by $\bm{I}^{{}^{\prime}}=\bm{I}+\bm{R}$.
@@ 3.2 Dynamic Radial Embedding
$\bm{D}_{\theta}\in\mathbb{R}^{H\times W}$ denotes the distribution matrix of polar coordinate and ${arctan}$ denotes the arctangent function. $\bm{D}_{x},\bm{D}_{y}\in\mathbb{R}^{H\times W}$ represents position matrices of the x-axis and the y-axis based on the Cartesian coordinate. Then, we use a polar mask to generate the offsets in each sector part by convolution layer. Given the ${N}_{1}$ as the number of azimuths in the polar coordinate system, we divided the full image into ${N}_{1}$ parts and generated the corresponding mask matrix. We combine the masking tensor with the convolution layer and capture the offsets from the directions of each azimuth.
$\Delta{p}_{i}=Conv_{3\times 3}(\bm{X},\bm{D}_{\theta\_{i}})\quad i=\{1,2,...,N_{1}\}$ (3)
@@ 3.3 Radial Strip Attention Solver
Swin transformer relative positional encoding contains the position bias $\bm{B}\in\mathbb{R}^{{M^{2}}\times{M^{2}}}$, M denotes the number of patches in a window, then the attention module computing is formulated as:
@@
where $\theta_{i}$ and $\phi_{i}$ denote the angular of $i$-th token, and the $\theta_{max}$ represents the half field of tensor. Then, the tokens $i$ and $j$ has the relative position $(\Delta\theta,\Delta\phi)$ can be given as:
@@
Then, the two position bias tensors $\bm{B}_{\theta}$ and $\bm{B}_{\phi}$ can be defined as:
@@
Corresponding to the polar coordinate-based patch embedding and self-attention, we design a polar azimuth patch merging. The neighboring patches are merged between radials to fit the strip window shapes.
@@ 4.2 Experimental Results
Table 6: Ablation study of DRE on GoPro [27] dateset.
@@
Table 7: Ablation study of RSAS on GoPro [27] dataset.

### v3

@@ Abstract
Exploring motion information is important for the motion deblurring task. Recently window-based transformer approaches have achieved decent performance in image deblurring. Note that the motion causing blurry results is usually composed of translation and rotation movements and the window-shift operation in the Cartesian coordinate system by the window-based transformer approaches only directly explores translation motion in orthogonal directions. Thus, these methods have the limitation of modeling the rotation part. To alleviate this problem, we introduce the polar coordinate-based transformer, which has the angles and distance to explore rotation motion and translation information together. In this paper, we propose a Radial Strip Transformer (RST), which is a transformer-based architecture that restores the blur images in a polar coordinate system instead of a Cartesian one. RST contains a dynamic radial embedding module (DRE) to extract the shallow feature by a radial deformable convolution. We design a polar mask layer to generate the offsets for the deformable convolution, which can reshape the convolution kernel along the radius to better capture the rotation motion information. Furthermore, we proposed a radial strip attention solver (RSAS) as deep feature extraction, where the relationship between windows is organized by azimuth and radius. This attention module contains radial strip windows to reweight image features in the polar coordinate, which preserves more useful information in rotation and translation motion together for better recovering the sharp images. Experimental results on six synthetic and real-world datasets prove that our method performs better than the SOTA methods for the image deblurring task. The code are available in the https://github.com/Calvin11311/RST.
@@ 1 Introduction
Image deblurring aims to recover a sharp image from the blurred one. This is a challenging problem as both blur and latent clear images are unknown. This problem has attracted many researchers from academia and industry due to the rapid development of kinds of smartphones. Significant progress has been made due to kinds of deep convolution neural networks (CNNs) [41, 26, 12]. The key success of these methods is because of the capability to model the local blur pattern by a convolution kernel and the diversity design of network architecture (e.g., the multi-scale [41, 43], multi-stage [26, 35] generative adversarial learning [17, 18], and latent space kernel estimation [12]). Since the limited receptive field and the spatially invariant of the convolution kernels [40] can not fully explore non-local contextual information that is vital for image deblurring, a deeper and larger network may relieve this problem. However, simply increasing the depth or capacity of deep networks does not always lead to better performance [44, 28].
@@
Figure 1: (a) Illustration of the motion field from the kernel estimation method [2], the red curves describe the motion state which causes the blurry images. (b) The demonstration of the motion trajectory causing blurry results can be composed of translation and rotation parts, as the blue curve consists of the red rotation motion and the yellow translation motion. (c) Illustrate the token relationship modeled under the polar coordinate system, where the $\theta_{ab}$ can represent the angular relative position between tokens (pixels) A and B.
@@
To overcome this problem, several approaches [40, 36, 16] explore the vision Transformer [10] to solve image deblurring and achieve decent performance on kinds of benchmark datasets. However, the scaled dot-product attention in the Transformer is time-consuming. Some methods adopt the rectangle window [6] or the strip windows [36] to compute self-attention on lower computation complexity. Since the restricted spatial extent of self-attention results in limitations for capturing long-range pixel relationships, the window shift operation is necessary to expand the receptive field. The shift operation under the Cartesian coordinate system captures the relationship between windows in orthogonal directions. However, in Figure 1 (a), we can clearly find that the blur kernel in real-world blurry images is an irregular curve indicating that the motion causing blurry images is complex. Furthermore, the motion can usually be decomposed into rotation and translation motion. Deng et al. [30] formulate the decomposition as $RT(o,\theta,t,x)=R(\theta)(x-o)+o+t$, where t and $\theta$ represent the translation vector and the rotation angle, respectively, $o$ and $x$ are the coordinates in the blurry image and the $R(\theta)$ is a geometric rotation matrix. Based on the above theory, as illustrated in Figure 1 (b), the trajectory of motion can consist of the rotation (the red curve) and the translation (the yellow line). We note that attention operations capturing both translation and rotation motion information together can enhance the ability to restore clearer images.
@@
However, shift operation under the Cartesian coordinate system (e.g., Stripformer [36]) only considers the cross-window connection in orthogonal directions, which explicitly models translation motion, while implicitly modeling rotation motion, leading to the limited capability to model the rotation information for restoring sharp images. In the polar coordinate system, we can use angles and distance to demonstrate the relative position, where the translation and the rotation can be considered together. So it is more suitable to develop an effective and efficient polar system-based transformer architecture to explore the motion information of blurry images better. In this paper, we propose a Radial-Strip Transformer (RST), exploiting two polar coordinate system-based modules: the dynamic radial embedding (DRE) and radial strip attention solver (RSAS), to capture the rotation and the translation features of motion efficiently. The DRE utilizes a polar mask to construct the shallow feature extractor, which is combined with the deformable convolution to capture the motion information of blur on the polar way. In detail, the rotation information can be captured by angles and the translation part is modeled by distance.
@@
Since features extracted from DRE are not fully utilized by attention operations with the Cartesian coordinate system, directly using window-based transformer architecture (e.g., Swin Transformer [24]) will not result in a good performance of image deblurring. Inspired by the DarSwin [1], which proposes a polar transformer under a wide-angle lens, we design an efficient radial strip attention solver (RSAS), which contains strip windows along the radius. Sampling operations are needed in sector window partitioning but can result in the loss of details, which is important for image deblurring. So we cancel the sampling operation and set the strip window, which is beneficial for preserving the useful detailed information for clear image restoration. Combined with the two designs above, we proposed our RST with an asymmetric encoder-decoder architecture, which only uses the RSAS in the decoder part. It is more efficient to solve the deblurring task because the shallow layers in attention may dilute the features of blur pattern [16]. The experimental results support that the method performs a favorable deblurring result against SOTA methods. The title “Spread your wings" gives us the imagination demonstrating a process of exploring more directions for capturing the motion information more fully under the polar coordinates.
@@
    We propose an efficient polar coordinate system-based embedding module DRE for extracting rotation and translation motion information of shallow features. It transforms the deformable convolution layer in each radial direction and preserves the information in rotation and translation.
@@ 2.1 Deep CNN-based approaches
Recently, the development of deep CNN methods [4, 7, 13] has achieved great performance in the image deblurring task. Nah et al. [26] design a muti-scale network to estimate the blur pattern from a blurry image directly. To make the multi-scale architecture extract the blur pattern more efficiently, Gao et al. [13] propose a parameter selective sharing method. [34] utilized the CNNs to predict the probabilistic distribution of the blurring kernel at the patch level. Moreover, the multi-patch architecture is another type to improve the performance of deblurring. Zhang et al. [43] perform a hierarchical deblurring model based on the multi-patch strategy, which estimates the feature of the blur pattern stage by stage. The cross-stage strategy is utilized by Zamir et al. [41], which fuses the different stage features to better restore the sharp image.
The above methods only directly use the fixed convolution kernel to estimate the blur pattern, which ignores the important movement prior of the blurry images. Some methods consider the direction of degradation, and WDNet [23] designs a dual-branch network for demoiréing, which contains a DPM block to capture the texture from eight directions. DSC [14] utilizes the attention mechanism in a spatial recurrent neural network for locating shadows from each direction. Fang et al. [11] try to estimate the non-uniform blur kernel in latent space and use it to reconstruct the sharp image. Due to the convolution operation being a spatial invariant and limited receptive field [40], CNN methods do not efficiently capture the global information for restoring blurry images.
@@ 2.2 Vision Transformer-based approaches
Transformer architecture achieves a significant performance of modeling global contexts with self-attention in high-level tasks (e.g., moving object segmentation [29], object detection [17], and text recognition [19]), and extend to low-level tasks(image deblurring [16, 40], super-resolution [21], and image denoising [3, 38]). But, self-attention brings a problem of huge computational cost. To solve it, SwinIR [21] divides the image by 8$\times$8 windows and executes an attention operation in each window. Restormer [40] designs across channel dimensions model to reduce the cost. The spatial information is important for the image restoration task, which is not fully developed. CAT [6] aggregates the features across rectangle windows with axial-shift operation, which can achieve a large receptive field and boost the interactions between windows with linear complexity. Stripformer [36] proposed an intra-strip and inter-strip attention mechanism, which decomposes motion information into horizontal and vertical directions for estimating the blur pattern. However, the above methods designed based on the Cartesian coordinate system can only explore the motion information in orthogonal directions. Since translation and rotation are both important parts of object movement, the Cartesian coordinate has a limited ability to model the rotation. Different from vanilla transformers calculate the relationship between pixels on the Cartesian system, we propose an efficient polar coordinate transformer model to explore the spacial information of movement with translation and rotation.
@@ 3.1 Overall Pipeline
Given a blurry image $\bm{X}$ with a resolution of $H\times W$. RST first adopts a deformable convolution layer with a radial mask to extract the shallow feature $\bm{F}_{0}\in\mathbb{R}^{H\times W\times C}$, where $H,~{}W,~{}C$ denote the height, width, and number of channels, respectively. Then the shallow feature $\bm{F}_{0}$ is embedded in the L-level encoder-decoder networks where each level has a skip connection between the encoder and decoder. Following the pioneered work [16], we directly use the frequency domain-based feed-forward network to refine the features, which has the benefit of capturing useful context information for deblurring. The process can be formulated as:
@@
$\mathcal{F}(\cdot)$ denotes the fast Fourier transform and $\mathcal{F}^{-1}(\cdot)$ denotes the inverse one. $\bm{W}$ represents a learnable quantization matrix. $\bm{F}_{e}\in\mathbb{R}^{H\times W\times C}$ represents the output feature from FFN of encoder parts. We use the transpose convolution for downsampling and upsampling and add the skip connection between the encoder and decoder on each level. In the decoder, we design radial strip attention (as shown in Figure 2) to extract the blur pattern from the rotation and translation movement. To be specific, the feature $\bm{F}_{a}\in\mathbb{R}^{\frac{H}{M}\times\frac{W}{M}\times C}$ ($M$ represents the patch size) product by window partition layer. After that, the windows are assembled by an azimuth patch merging module with the angular relative positional encoding. The transformer block outputs the final restored feature $\bm{F}_{r}\in\mathbb{R}^{H\times W\times C}$, and using a $3\times 3$ convolution layer to generate the residual part $\bm{R}\in\mathbb{R}^{3\times H\times W}$. Finally, the restored image is generated by $\bm{X}^{{}^{\prime}}=\bm{X}+\bm{R}$.
@@ 3.2 Dynamic Radial Embedding
$\bm{D}_{\theta}\in\mathbb{R}^{H\times W}$ denotes the distribution matrix of polar coordinate and ${arctan}$ denotes the arctangent function. $\bm{D}_{x},\bm{D}_{y}\in\mathbb{R}^{H\times W}$ represents position matrices of the x-axis and the y-axis based on the Cartesian coordinate. Then, we use a polar mask to generate the offsets in each sector part by convolution layer. Given the ${N}_{1}$ as the number of azimuths in the polar coordinate system, we divided the full image into ${N}_{1}$ parts and generated the corresponding mask matrix. We combine the mask matrix with the convolution layer and capture the offsets from the directions of each azimuth.
$\Delta{p}_{i}=Conv_{3\times 3}(\bm{X}\cdot\bm{D}_{\theta\_{i}})\quad i=\{1,2,...,N_{1}\}$ (3)
@@ 3.3 Radial Strip Attention Solver
Swin transformer relative positional encoding contains the position bias $\bm{B}\in\mathbb{R}^{{M^{2}}\times{M^{2}}}$, $M$ denotes the number of patches in a window. The query, key, and value matrices $Q,K,V$ are from the projections matrices $P_{Q},P_{K},P_{V}$, shared across windows. Then the attention module computing is formulated as:
@@
where $\theta_{i}$ and $\phi_{i}$ denote the angular of $i$-th token, and the $\theta_{max}$ represents the half field of tensor. $N_{\phi}$ and $N_{r}$ denote the division number of angles and the length in the radial direction separately. Then, the tokens $i$ and $j$ has the relative position $(\Delta\theta,\Delta\phi)$ can be given as:
@@
The $N_{\theta}$ represents the cut numbers in the radial direction. Then, the two position bias tensors $\bm{B}_{\theta}$ and $\bm{B}_{\phi}$ can be defined as:
@@
Corresponding to the polar coordinate-based patch embedding and self-attention, we design a polar azimuth patch merging. The process of patch merging is similar to other window-based attention (e.g., Swin transformer [24]), the difference is that our patch merged along the azimuth due to the radial strip window.
@@ 4.2 Experimental Results
Table 6: Evaluations of DRE with different sector numbers on GoPro [26] dataset.
@@
Table 7: Quantitative evaluations of each module in proposed method on the GoPro [26] dataset. The number of sector is 4.
@@ 5 Conclusion
Limitation. Our method still has some shortcomings for future improvement. Although we designed the radial strip window to reduce the computational cost, which also leads the insufficient cross-window interactions. For the heavy blurry in overly complex real-world scenarios, our proposed method still has limited capacity, as shown in the failure cases in supplemental materials. Creating a dataset that can cover the various real-world might be a solution for improvements